OpenStudy (anonymous):

5. Using complete sentences, explain how to find the average rate of change for f(x) from x = 4 to x = 7. my equation is F(x) = 3 (x-3)^2 + 4 in vertex form and and y = 3x^2 - 18x +31. Any ideas how to do this would be super helpful and I'll give you a medal!

OpenStudy (anonymous):

y = 3x^2 - 18x +31 in standard form

OpenStudy (campbell_st):

use the slope formula... you need 2 points average rate of change is \[\frac{f(b) - f(a)}{b - a}\] so substitute the x values to get a corresponding f(x) then substitute into the formula

OpenStudy (anonymous):

How do I do that? Like I know I need a y correct? So do I solve for x plug that in the parabola formula and use the y and find the slope?

OpenStudy (campbell_st):

of so looking at the equation in vertex form find F(7) \[F(7) = 3(7 -3)^2 + 4 = 52\] and for x = 4 \[F(4) = 3(4 -3)^2 + 4 = 7\] so the average rate of change is \[\frac{52 - 7}{7 - 4} = \] same process for the standard form equation.

OpenStudy (campbell_st):

oops just realised they are the same equations... so all you need to do is simplify the fraction for the average rate of change...

OpenStudy (anonymous):

So you plug in the number for x, solve for y then you get that fraction?

OpenStudy (campbell_st):

you do... you actually get 2 points as shown above (7, 52) and (4, 7) so all you need to do for the average rate of change is find the slope of the interval joining the 2 points

OpenStudy (anonymous):

And how would I do that?

OpenStudy (campbell_st):

well you you go up to my 1st post you have (52 - 7)/(7 -4) = just evaluate it

OpenStudy (anonymous):

is it 15?...

OpenStudy (campbell_st):

thats it... the average rate of change.

OpenStudy (anonymous):

So the average rate of change is 15? I feel silly for even asking that question haha... Thank you :)

OpenStudy (campbell_st):

yes... its the average rate of change in x with respect to y... glad to help